OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (-8, 0), (1, 5)

OpenStudy (anonymous):

Before we input the equation of the line, since 2 points in the line are given, we can solve for the slope of the line...\[m=\frac{ (5-0) }{ (1-(-8)) }=\frac{ 5 }{ 9 }\]

OpenStudy (anonymous):

and from definition of slope of a line, other points satisfy the equation of a line considering a given point, either (-8,0) or (1,5)... other points (x,y)....\[m = \frac{ (y-y _{1}) }{ (x-x _{1}) }\]

OpenStudy (anonymous):

choosing (1,5) as the point and the slope m=5/9... we have...\[\frac{ 5 }{ 9 }=\frac{ (y-5) }{ (x-1) }\]

OpenStudy (anonymous):

5(x-1) = 9(y-5) ... by cross multiplication 5x-5 = 9y-45 ... by distributive property 5x - 9y - 5 + 45 = 0 ... by transposistion 5x - 9y + 40 = 0 ... the equation of the line

OpenStudy (anonymous):

to check, we can substitute the other point (-8,0)... 5 (-8) - 9 (0) + 40 = 0